Show that the statement, p: if a is a real number such that a3 + 4a =0, then a is 0″, is true by direct method?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let q and r are the statements given by q: a is a real number such that a^3+ 4a =0

r: a is 0.

let q be true then

a is a real number such that a^3+ 4a =0

a is a real number such that a(a^2+ 4) =0

a = 0

r is true

So, q is true and r is true,
so p is true.

Thanks